Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form and explain each step you took to get to your answer.
Answer by lwsshak3(11628)

Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form and explain each step you took to get to your answer.
**
point-slope formula: (y1-y2)=m(x1-x2)
find slope,m=∆y/∆x=change in y/change in x
using coordinates of given points,
m=(4-(-8))/(3-(-3))
=(4+8)/(3+3)
=12/6
m=2
apply point-slope formula using any one of the given points(3,4), and slope m=2
(y-4)=2(x-3)